OpenStudy (anonymous):

if a car travels 50 miles in 2 hours, how many hours will it take to travel 90 miles?

OpenStudy (amistre64):

90*2/50 or 50/(90*2) cant recall which

OpenStudy (asnaseer):

50 miles in 2 hours which is 120 minutes so 1 mile in 120/50 minutes = 12/5 minutes so 90 miles in 90*(12/5) minutes = 18*12 minutes = 216 minutes = 3 hours 36 minutes

OpenStudy (anonymous):

thankssssss (: